Need Help I Answer some but stuck.

  1. Determine the amount of sales (units) that would be necessary under

    Break-Even Sales Under Present and Proposed Conditions

    Darby Company, operating at full capacity, sold 70,200 units at a price of $57 per unit during the current year. Its income statement for the current year is as follows:

    Sales $4,001,400
    Cost of goods sold 1,976,000
    Gross profit $2,025,400
    Expenses:
    Selling expenses $988,000
    Administrative expenses 988,000
    Total expenses 1,976,000
    Income from operations $49,400

    The division of costs between fixed and variable is as follows:

    Variable Fixed
    Cost of goods sold 70% 30%
    Selling expenses 75% 25%
    Administrative expenses 50% 50%

    Management is considering a plant expansion program that will permit an increase of $342,000 in yearly sales. The expansion will increase fixed costs by $34,200, but will not affect the relationship between sales and Costs that vary in total dollar amount as the level of activity changes.variable costs.

    Required:

    1. Determine the total variable costs and the Costs that tend to remain the same in amount, regardless of variations in the level of activity.total fixed costs for the current year. Enter the final answers rounded to the nearest dollar.
    Total variable costs $2,618,200.00 $
    Total fixed costs $1,333,800.00 $
    1. Determine (a) the unit variable cost and (b) the The dollars available from each unit of sales to cover fixed costs and provide operating profits.unit contribution margin for the current year. Enter the final answers rounded to two decimal places.
    Unit variable cost $ 37.30
    Unit contribution margin $ 19.70
    1. Compute the break-even sales (units) for the current year. Enter the final answers rounded to the nearest whole number.
      units $
    2. Compute the break-even sales (units) under the proposed program for the following year. Enter the final answers rounded to the nearest whole number.
      units $
    3. Determine the amount of sales (units) that would be necessary under the proposed program to realize the $49,400 of income from operations that was earned in the current year. Enter the final answers rounded to the nearest whole number.
      units $
    4. Determine the maximum income from operations possible with the expanded plant. Enter the final answer rounded to the nearest dollar.
      $
    5. If the proposal is accepted and sales remain at the current level, what will the income or loss from operations be for the following year? Enter the final answer rounded to the nearest dollar.
      $ Income $
    • Income
    1. Based on the data given, would you recommend accepting the proposal?
    1. In favor of the proposal because of the reduction in break-even point.
    2. In favor of the proposal because of the possibility of increasing income from operations.
    3. In favor of the proposal because of the increase in break-even point.
    4. Reject the proposal because if future sales remain at the current level, the income from operations will increase.
    5. Reject the proposal because the sales necessary to maintain the current income from operations would be below the current year sales.

    Choose the correct answer.
    b

    • a
    • b
    • c
    • d
    • e

    Feedback

    1. Multiply the percentages for fixed and variable costs by each cost.

    2. a. Divide the total variable costs by number of units.

    2. b. Sales price per unit minus variable costs per unit equals contribution margin per unit.

    3. Fixed costs divided by unit contribution margin equals break-even point.

    4. Fixed costs under the proposed program divided by contribution margin equals new break-even point.

    5. (Fixed costs + Target profit) divided by unit contribution margin equals sales units.

    6. Determine the increase in units by dividing the sales increase by the price per unit. Add the additional revenue and additional fixed costs when calculating:

    Sales minus fixed and variable costs equals income from operations.

    7. Subtract the additional fixed costs from the operating income.

    8. Consider the break-even point and the sales needed for the proposed level.

    Learning Objective 2, Learning Objective 3.

Loading item

There was an error loading this item. If this continues to occur, please contact Technical Support.

Check My Work

  • Previous
  • Next
  • 100% Correct
  • Partially Correct
  • Incorrect
  • Needs Instructor Grading

Basic Calculatorclose

0

UseEntBSBSpCEHomCEnd

789+

456-

123*

0.=/
Answer #1

Solution

Solution

Darby Company

Determination of total variable costs and the costs that tend to remain the same in amount regardless of variations in the level of activity:

Total variable costs –

Cost of goods sold      70% x 1,976,000 = $1,383,200

Selling expenses         75% x 988,000 = $741,000

Administrative expenses 50% x 988,000 = $494,000

Total variable expenses = $2,618,200

Total fixed costs –

Cost of goods sold      30% x 1,976,000 = $592,800

Selling expenses         25% x 988,000 = $247,000

Administrative expenses 50% x 988,000 = $494,000

Total fixed cost = $1,333,800

Unit Variable cost = total variable cost/units sold

= $2,618,200/70,200 units

= $37.30 per unit

Unit Contribution margin = unit selling price – unit variable cost

= $57 – 37.30 = $19.70

Determination of the break-even unit sales for current year:

Break-even unit sales = total fixed cost/contribution margin per unit

= $1,333,800/$19.70 = 67,706 units

Determination of the break-even unit sales under the proposed program for the following year:

Proposed program –

Increase in yearly sales = $342,000

Increase in fixed cost = 34,200

Revised sales = $4,001,400 + $342,000 = $4,343,400

Percent increase in sales = 342,000/4,001,400 = 8.55%

Increased variable cost = 2,618,200 + 8.55% x 2,618,200 = $2,842,000

Increased fixed cost = $1,333,800 +34,200 = $1,368,000

Contribution margin = sales – variable cost

= 4,343,400 – 2,842,000 = $1,501,400

Sales units = $4,343,400/$57 = 76,200 units

Contribution margin per unit = $1,501,400/76,200 units = $19.70

Break-even sales units under the proposed program = $1,368,000/$19.70 = 69,442 units

Determination of the amount of sales needed to realize $49,400 income from operations under the proposed program:

Desired sales = (target income + fixed cost)/contribution margin

Target income = $49,400

Fixed cost –

Original fixed cost      $1,333,800

Add: increase              $34,200

Total fixed cost           $1,368,000

Contribution margin = sales – variable cost

Sales price = $57

Variable cost per unit = $37.30

Contribution margin per unit = $19.70

Desired sales = (49,400 + 1,368,000)/$19.70 = 71,949 units

Determination of maximum income possible from expanded plant:

Revised Sales             $4,343,400

Less: Variable costs    $2,842,000

Contribution margin   $1,501,400

Less: Fixed cost          $1,368,000

Net operating income $133,400

If proposal accepted and sales remain at current level, determination of net income:

Sales                           $4,001,400

Variable cost               $2,618,200

Contribution margin   $1,383,200

Fixed cost                   $1,368,000

Net operating income $15,200

Based on the given data, the proposed program could be accepted –

In favour of the proposal because of the possibility of increasing income from operations.

The income from operations under the proposed program would be $133,400. Hence, the proposed program would increase income from operations by $84,000 ($133,400 – 49,400).